User name J. Singh
Log entry time 21:46:40 on July12,2003
Entry number 103785
keyword=Laser Alignment + Settings
Patricia + JP aligned the lasers starting @ 1730. Finished at about 20:30. it went "well...wow."
started "spinup" @ 0 deg @ 2030. note that the pumping chamber temps seem higher than usual. laser hitting the RTD? (green = RTD6, red = RTD7, blue = oven RTD).
laser settings (should be optimized...):
Long #4: 38.0A / 14.3C ON
Long #5: 37.5A / 17.9C OFF
Long #7: 40.0A / 08.3C ON
Long #6: 38.1A / 10.0C ON
Trans #1: 36.0A / 15.3C
Trans #2: 41.4A / 07.2C
Trans #3: 38.1A / 10.8C